### v1

@@ 2.2.3 Spatiotemporal Patterns Division
Classic traffic theory forms the foundation by classifying traffic into free flow and congested flow, where congested flow encompasses a variety of spatiotemporal traffic patterns. Free flow is characterized by high vehicle speeds and low densities, with minimal interaction between vehicles. Congested flow, on the other hand, includes a range of traffic states such as stop-and-go waves, wide-moving jams, and localized clusters. IDM, through its nonlinear acceleration framework, can reproduce several distinct traffic patterns, including pinned localized clusters (PLC), moving localized clusters (MLC), triggered stop-and-go waves (TSG), oscillatory congested traffic (OCT), and homogeneous congested traffic (HCT) (Treiber et al., 2010). These patterns arise continuously, with smooth transitions between states, as shown in Fig. 3. This flexibility and ability to model traffic instabilities have established IDM as a benchmark model for simulating traffic dynamics within the two-phase framework.
Figure 3: Dynamic phase diagram of on-ramp-induced congested traffic patterns simulated by IDM from Treiber et al. (2010). The abbreviations denote free traffic (FT), pinned localized cluster (PLC), moving localized cluster (MLC), homogeneous congested traffic (HCT), oscillatory congested traffic (OCT), and triggered stop-and-go (TSG) pattern.
@@ 3   Critical Discussion of the IDM and Its Extensions
{tblr}
width = colspec = Q[150]Q[160]Q[350], cells = m, hline1,10 = -0.08em, hline2 = -, rowodd = gray9, Model & Improvements / Features Use Cases
@@ 3.2.1 Deterministic IDM Extensions
For example, Treiber et al. (2006) emphasized the importance of incorporating reaction time and human factors. While IDM does not directly account for reaction time, it can be extended to include factors such as reaction delays, estimation errors, and multi-vehicle anticipation. Although these factors can undermine traffic stability, they may be compensated by introducing spatial and temporal expectation mechanisms. This explains why IDM and its deterministic extensions, such as the Human Driver Model (HDM), effectively reproduce empirical traffic dynamics despite their simplicity. It further stabilizes traffic flow by increasing oscillation wavelength and reducing the gradient of stop-and-go waves, aligning better with empirical observations. Similarly, Kesting et al. (2008) extended IDM for simulating adaptive cruise control systems, creating the IDM-ACC, which captures smoother and more stable driving patterns characteristic of ACC-equipped vehicles. Studies such as Li et al. (2017a, b) demonstrated IDM-ACC’s ability to evaluate the impact of CACC systems on highway safety and efficiency, showing its robust applicability in automated driving. Despite differences in model components, HDM and IDM-ACC often produce similar patterns due to their shared foundation in IDM’s "intelligent braking strategy" (see subsections 2.1.2 and 2.2.1).
Other deterministic extensions further refine IDM to address specific weaknesses. For instance, Improved IDM (I-IDM) proposed by Treiber and Kesting (2013b) eliminates exaggerated deceleration at speeds exceeding the desired speed $v_{0}$, ensuring equilibrium gaps align with empirical observations. Sigmoid-IDM (Chen et al., 2024b) incorporates a sigmoid function to better handle traffic oscillations, while Task-Difficulty IDM (TD-IDM) (Saifuzzaman and Zheng, 2014) accounts for task-difficulty considerations to model human factors in heterogeneous environments.
@@ 3.2.2 Stochastic IDM Extensions
Jiang et al. (2014) introduced the two-dimensional IDM (2D-IDM), which revealed the concave growth pattern of traffic oscillations through time-varying headway analysis. Building on this, Xiong and Jiang (2022) utilized 2D-IDM to establish a dynamic programming model for recommended speed at isolated signalized intersections in connected environments, characterizing random human driving behavior. Further advancements were made by Tian et al. (2016b), who proposed the Improved 2D-IDM (I2D-IDM). This model successfully simulated synchronized flow and traffic oscillations by incorporating defensive driving behavior at high speeds, thus improving realism in high-speed scenarios.
To address traffic oscillations, Treiber and Kesting (2018) introduced IDM with white noise and IDM with action points, showing that even basic stochastic elements can effectively reproduce observed fluctuation patterns. Similarly, Tian et al. (2019) proposed the Stochastic Speed Adaptation Model (SSAM), which utilized the function formulation of desired spacing of IDM (see Equation (2)) and captures critical traffic instability characteristics such as speed adaptation and spacing indifference. By doing so, SSAM successfully reproduces the concave growth patterns observed in empirical data.
Other stochastic extensions have focused on modeling inter-driver and intra-driver variability. For instance, Lindorfer et al. (2018) presented the Enhanced HDM, which incorporates stochastic Wiener processes to model driving errors, variable reaction times, and driver distractions. Zhang et al. (2022) proposed a time-varying IDM that emphasizes both inter-driver and intra-driver heterogeneity, enhancing the model’s ability to simulate real-world variability in human driving behavior.
@@ 3.2.3 IDM Extensions for Automated Driving
In addition to the characterization of driving behavior, IDM has also been extended to various aspects of autonomous driving modeling, such as obstacle avoidance, vehicle merging, ecological driving, network attacks, and information error modeling. Sharath and Velaga (2020) proposed AV-IDM for two-dimensional motion planning, specifically in mixed traffic scenarios involving obstacle avoidance in which the surrounding obstacles like vehicles and road boundaries are treated as stimuli. Chen et al. (2024a) developed a model focusing on freeway merging areas, redefining the desired safety distance in IDM to enhance safety during merging under AV environments. Ward et al. (2017) proposed a probabilistic extension of IDM for AVs, incorporating normal distribution noise to improve interaction-aware planning in merge scenarios. Zhou et al. (2017) introduced the Multi-AV cooperative (MA-C) IDM, integrating cooperative components to assess freeway merging under various AV ratios. Zhou et al. (2023a) introduced the EIDM for cooperative control in mixed traffic, combining multiple vehicle state information to mitigate the negative impacts of HVs. As to eco-driving of AV, Dong et al. (2024) proposed a Flexible Eco-Cruising Strategy for CAVs, employing a hierarchical control framework to optimize driving lane planning and speed, thus reducing driving costs in various traffic conditions. When addressing cybersecurity and information errors,Wang et al. (2020a) developed the CIDM to model CAV behavior under cyberattacks, utilizing dynamic communication topologies to analyze the impact of manipulated information. Berkhahn et al. (2022) introduced IDMrm, incorporating random misperception (rm) into IDM to model traffic dynamics at intersections with potential perception errors. Li et al. (2024b)presented the Dynamic Cooperative IDM (DC-IDM), addressing information errors and their impact on CAVs. Liu et al. (2022a) proposed a microscopic model for mixed traffic flow unsignalized intersections for autonomous driving, which incorporates the Ornstein–Uhlenbeck process into IDM to model the random perception errors of vehicles at intersections. Luo et al. (2024) proposed the VC-IDM, integrating a generalized force model to analyze platoon behavior and defense mechanisms under cyberattack scenarios.
@@
Summary: The IDM and its extensions owe much of their success to the foundational "intelligent braking strategy," which ensures collision-free behavior when physically possible Treiber and Kesting (2013b). This feature not only enhances IDM’s safety and robustness but also serves as the unifying principle across all deterministic and stochastic variants, making IDM a reliable and flexible framework for both research and practical applications. In deterministic extensions, models like IDM-ACC and HDM refine IDM’s ability to reproduce empirical traffic dynamics, improve stability, and adapt to automated driving scenarios. Meanwhile, stochastic extensions, such as 2D-IDM and SSAM, introduce variability to better capture real-world phenomena like driver heterogeneity, traffic oscillations, and random behavior, thereby enhancing IDM’s realism in diverse traffic conditions.
@@ 3.3.1 Reinforcement Learning-Based Models
Reinforcement learning (RL)-based extensions of IDM aim to improve decision-making and vehicle control by integrating data-driven adaptability with IDM’s physically interpretable structure. For instance, Brito et al. (2022) extended IDM using deep reinforcement learning to predict and identify leading vehicles, estimate motion plans, compute control commands (e.g., acceleration), and optimize trajectories in dense traffic, thereby enhancing safety and feasibility. Similarly, Cao et al. (2022b) developed a decision framework for RL, incorporating IDM as a lower performance bound, where autonomous vehicles (AVs) only adopt RL policies if they outperform IDM.
Additionally, Bai et al. (2024) combined Twin Delayed Deep Deterministic Policy Gradient (TD3)-based RL with IDM to achieve improved longitudinal control, safety, and efficiency. Kerbel et al. (2024) incorporated IDM’s expected acceleration as a state variable in shared learning to enhance fleet-wide fuel economy and adaptability. Furthermore, Hart et al. (2024) proposed an RL model with reward functions grounded in IDM’s assumptions, using parameters such as the desired time gap, maximum and comfortable accelerations/decelerations, and desired speed to guide learning.
@@ 3.3.2 Machine and Deep Learning-Based Models
The integration of IDM with machine learning (ML) and deep learning (DL) methods has enabled advanced modeling of vehicle behavior by blending data-driven insights with the established physical frameworks of IDM. For example, Chen et al. (2020) developed an LSTD-IDM model that combines long-term and short-term driving characteristics by merging IDM with longitudinal driving data. Mo et al. (2021) proposed a physics-informed deep learning framework (PIDL-CF), blending IDM with deep learning to predict driver acceleration under various conditions.
Hybrid models have also extended IDM with probabilistic and transformer-based approaches. Bhattacharyya et al. (2022) introduced a hybrid particle-filtering method combined with collaborative IDM extensions. Geng et al. (2023) presented the Transformer-IDM (PIT-IDM) for enhanced vehicle trajectory prediction by integrating IDM with deep learning techniques. Similarly, Kang et al. (2023) proposed PGM-IDM, combining trajectory-based embeddings with probabilistic graphical models to improve vehicle-following responses.
Further advancements include Liu et al. (2023a), which introduced QRPIDL-CF, fusing quantile regression and deep learning with IDM’s physical insights. Tang et al. (2024) proposed a grounded relational reasoning approach that incorporates IDM’s expected distance into the reward function of a Markov decision process, while Zhang and Sun (2024) developed a memory-augmented Bayesian calibration (MA-IDM) to address parameter uncertainties and behavioral differences in IDM, significantly enhancing its predictive accuracy.
@@ 3.4.1 Deterministic Model Comparisons
Many studies have compared their proposed models with IDM, but relatively few have conducted comprehensive and unified evaluations that fairly compare all models, including IDM. One such study is Punzo et al. (2021), where deterministic models, including IDM, Gipps’ model, and the full velocity difference model (FVDM) with constant-time headway (CTH) and SIGMOID spacing policies, were systematically evaluated across multiple datasets and calibration settings. IDM demonstrated strong and consistent calibration performance, achieving low errors in speed and acceleration while performing well in spacing accuracy. Although Gipps’ model slightly outperformed IDM in reproducing spacing for both human-driven and autonomous vehicle datasets, IDM maintained a more balanced performance across all three measures: spacing, speed, and acceleration. The robustness of IDM across diverse datasets and calibration settings, especially when using Pareto-efficient objective functions such as NRMSE(s,v,a), underscores its versatility as a reliable model for CF dynamics.
@@ 4.1 Applications in Traffic Stability Analysis
A significant focus has been on understanding traffic stability through IDM. Kesting and Treiber (2008) initiated this exploration by analyzing the influence of reaction time and speed adaptation on traffic stability, using IDM to reveal various instability mechanisms. Building on this, Jin (2016) demonstrated the equivalence between continuum and CF models, illustrating that higher-order IDM can be transformed into continuous models with consistent stability conditions. Monteil et al. (2014) advanced the discussion by examining string stability in mixed traffic flows and the impact of multiple anticipations in cooperative CF models. Montanino et al. (2021) further delved into string stability in mixed and heterogeneous traffic flows, incorporating uncertainties in model parameters. They later proposed a unified modeling framework (Montanino and Punzo, 2021) to assess stability, considering driver and vehicle heterogeneity and using IDM to verify this framework. Bouadi et al. (2022) extended these concepts into the realm of stochastic models, investigating string stability in stochastic CF models using IDM with stochastic variants and the generalized Lyapunov equation to establish stability conditions.
@@ 4.2 Applications in Calibration Framework
For the development of calibration frameworks, building on this focus on error, Jin et al. (2014) explored error accumulation in CF models calibrated with vehicle trajectory data, deriving a dynamic error model based on acceleration. Long et al. (2024) introduced a bi-scale CF model calibration using corridor-level trajectories, highlighting macro measurement error accumulation. To address the validation aspect, Treiber and Kesting (2012) developed a quantitative method for calibrating and validating traffic flow models based on the spatiotemporal evolution of congested traffic patterns. In terms of parameter importance and simplification, Punzo et al. (2015) conducted sensitivity analysis on driving behavior models, finding leader trajectory more important than parameters, and proposed a simplified model with faster convergence. For setting robust calibration objectives, Punzo et al. (2021) proposed robust guidelines for selecting MoP and GoF for calibrating CF dynamics. They used Pareto efficiency and indifference curves for function comparison. Focusing on data integrity, Sharma et al. (2019) investigated the impact of vehicular trajectory completeness on the calibration and validation of CF models. Finally, addressing behavioral pattern recognition, Zheng et al. (2023) proposed a multi-objective calibration framework to capture behavioral patterns of autonomously-driven vehicles, validating with empirical observations of ACC vehicle dynamics. Zhou et al. (2025) established the calibration framework of stochastic CF models and revealed the essential differences between deterministic and stochastic CF model calibration using various stochastic IDM variants.
@@
As for the development of the optimization algorithms for calibration, Rahman et al. (2015) proposed a stochastic calibration method using Bayesian estimation and MCMC simulation. Li et al. (2016) developed a global optimization algorithm combining direct and gradient searches for CF models. Zhong et al. (2016) introduced a calibration method using the Cross-Entropy Method and sensitivity analysis under noisy data. Fard and Mohaymany (2019) suggested a copula-based algorithm to explore complex parameter interactions in traffic models. Wang et al. (2024c) created an integrated macro-micro calibration framework using enhanced optimization and deep learning.
@@
In addition to the above theme, calibration is crucial for model evaluation. Zhu et al. (2018) investigated the applicability of CF models for Chinese drivers, highlighting IDM’s effectiveness and portability on Shanghai highways. Campi et al. (2024) assessed various models in roundabout scenarios, with IDM excelling in capturing CAV behavior. Zhou et al. (2023b) used four classical emission and fuel consumption models and found IDM effective at the vehicle pair level for energy consumption and emissions predictions. However, at the platoon level, predicted results were almost constant, contrasting with experimental observations.
@@
Based on calibrated IDM, different behavior patterns and vehicle types can be extracted. Ossen and Hoogendoorn (2011) used data collected by helicopter to calibrate the IDM to identify differences in following behaviors between different driver types, showing significant heterogeneity. Makridis et al. (2020) estimated the response time of vehicles equipped with ACC by calibrating IDM parameters and found that it was comparable to that of human drivers. Hu et al. (2023) calibrated IDM parameters to distinguish the empirical following behaviors of AVs and HVs in Waymo data. Li et al. (2023a) proposed an AV identification algorithm based on IDM calibration results, that is, using error results to distinguish AVs from HVs.
@@
Summary: IDM has played a pivotal role in the development of calibration frameworks, supporting advances in model performance metrics, optimization algorithms, and behavioral feature extraction. From error modeling to cross-scenario validation, IDM has been widely applied due to its analytical tractability and interpretability. However, most calibration efforts still focus on deterministic model variants, while the calibration of stochastic IDM-family models remains notably underdeveloped due to challenges in parameter identifiability, noise disentanglement, and the lack of robust likelihood functions (Punzo et al., 2021). This limits their empirical applicability despite their theoretical promise. Moving forward, calibration should be re-conceptualized not merely as parameter fitting but as a behavioral inference task, requiring methods that can accommodate uncertainty, heterogeneity, and model adaptivity across varying traffic regimes.

### v2

@@ 2.2.3 Spatiotemporal Patterns Division
Classic traffic theory forms the foundation by classifying traffic into free flow and congested flow, where congested flow encompasses a variety of spatiotemporal traffic patterns. Free flow is characterized by high vehicle speeds and low densities, with minimal interaction between vehicles. Congested flow, on the other hand, includes a range of traffic states such as stop-and-go waves, wide-moving jams, and localized clusters. IDM, through its nonlinear acceleration framework, can reproduce several distinct traffic patterns, including pinned localized clusters (PLC), moving localized clusters (MLC), triggered stop-and-go waves (TSG), oscillatory congested traffic (OCT), and homogeneous congested traffic (HCT) (treiberThreephaseTrafficTheoryTwophase2010). These patterns arise continuously, with smooth transitions between states, as shown in Fig. 3. This flexibility and ability to model traffic instabilities have established IDM as a benchmark model for simulating traffic dynamics within the two-phase framework.
Figure 3: Dynamic phase diagram of on-ramp-induced congested traffic patterns simulated by IDM from treiberThreephaseTrafficTheoryTwophase2010. The abbreviations denote free traffic (FT), pinned localized cluster (PLC), moving localized cluster (MLC), homogeneous congested traffic (HCT), oscillatory congested traffic (OCT), and triggered stop-and-go (TSG) pattern.
@@ 3  Critical Discussion of the IDM and Its Extensions
Model Improvements / Features Use Cases
@@ 3.2.1 Deterministic IDM Extensions
For example, treiber_delays_2006 emphasized the importance of incorporating reaction time and human factors. While IDM does not directly account for reaction time, it can be extended to include factors such as reaction delays, estimation errors, and multi-vehicle anticipation. Although these factors can undermine traffic stability, they may be compensated by introducing spatial and temporal expectation mechanisms. This explains why IDM and its deterministic extensions, such as the Human Driver Model (HDM), effectively reproduce empirical traffic dynamics despite their simplicity. It further stabilizes traffic flow by increasing oscillation wavelength and reducing the gradient of stop-and-go waves, aligning better with empirical observations. Similarly, kestingAdaptiveCruiseControlDesign2008 extended IDM for simulating adaptive cruise control systems, creating the IDM-ACC, which captures smoother and more stable driving patterns characteristic of ACC-equipped vehicles. Studies such as li_evaluation_2017; li_evaluating_2017 demonstrated IDM-ACC’s ability to evaluate the impact of CACC systems on highway safety and efficiency, showing its robust applicability in automated driving. Despite differences in model components, HDM and IDM-ACC often produce similar patterns due to their shared foundation in IDM’s "intelligent braking strategy" (see subsections 2.1.2 and 2.2.1).
Other deterministic extensions further refine IDM to address specific weaknesses. For instance, Improved IDM (I-IDM) proposed by treiberTrafficFlowDynamics2013 eliminates exaggerated deceleration at speeds exceeding the desired speed $v_{0}$, ensuring equilibrium gaps align with empirical observations. Sigmoid-IDM (chen_sigmoid-based_2024) incorporates a sigmoid function to better handle traffic oscillations, while Task-Difficulty IDM (TD-IDM) (saifuzzamanIncorporatingHumanFactorsCarFollowingModels2014) accounts for task-difficulty considerations to model human factors in heterogeneous environments.
@@ 3.2.2 Stochastic IDM Extensions
jiangTrafficExperimentRevealsNature2014 introduced the two-dimensional IDM (2D-IDM), which revealed the concave growth pattern of traffic oscillations through time-varying headway analysis. Building on this, xiong_speed_2022 utilized 2D-IDM to establish a dynamic programming model for recommended speed at isolated signalized intersections in connected environments, characterizing random human driving behavior. Further advancements were made by tianImproved2DIntelligentDriver2016, who proposed the Improved 2D-IDM (I2D-IDM). This model successfully simulated synchronized flow and traffic oscillations by incorporating defensive driving behavior at high speeds, thus improving realism in high-speed scenarios.
To address traffic oscillations, treiberIntelligentDriverModelStochasticity2018 introduced IDM with white noise and IDM with action points, showing that even basic stochastic elements can effectively reproduce observed fluctuation patterns. Similarly, tianRoleSpeedAdaptationSpacing2019 proposed the Stochastic Speed Adaptation Model (SSAM), which utilized the function formulation of desired spacing of IDM (see Equation (2)) and captures critical traffic instability characteristics such as speed adaptation and spacing indifference. By doing so, SSAM successfully reproduces the concave growth patterns observed in empirical data.
Other stochastic extensions have focused on modeling inter-driver and intra-driver variability. For instance, lindorfer_modeling_2018 presented the Enhanced HDM, which incorporates stochastic Wiener processes to model driving errors, variable reaction times, and driver distractions. zhangGenerativeCarFollowingModelConditioned2022c proposed a time-varying IDM that emphasizes both inter-driver and intra-driver heterogeneity, enhancing the model’s ability to simulate real-world variability in human driving behavior.
@@ 3.2.3 IDM Extensions for Automated Driving
In addition to the characterization of driving behavior, IDM has also been extended to various aspects of autonomous driving modeling, such as obstacle avoidance, vehicle merging, ecological driving, network attacks, and information error modeling. sharath_enhanced_2020 proposed AV-IDM for two-dimensional motion planning, specifically in mixed traffic scenarios involving obstacle avoidance in which the surrounding obstacles like vehicles and road boundaries are treated as stimuli. chen_safety_2024 developed a model focusing on freeway merging areas, redefining the desired safety distance in IDM to enhance safety during merging under AV environments. ward_probabilistic_2017 proposed a probabilistic extension of IDM for AVs, incorporating normal distribution noise to improve interaction-aware planning in merge scenarios. zhou_impact_2017 introduced the Multi-AV cooperative (MA-C) IDM, integrating cooperative components to assess freeway merging under various AV ratios. zhou_cooperative_2023 introduced the EIDM for cooperative control in mixed traffic, combining multiple vehicle state information to mitigate the negative impacts of HVs. As to eco-driving of AV, dongFlexibleEcoCruisingStrategyConnected2024a proposed a Flexible Eco-Cruising Strategy for CAVs, employing a hierarchical control framework to optimize driving lane planning and speed, thus reducing driving costs in various traffic conditions. When addressing cybersecurity and information errors,wang_modeling_2020 developed the CIDM to model CAV behavior under cyberattacks, utilizing dynamic communication topologies to analyze the impact of manipulated information. berkhahn_traffic_2022 introduced IDMrm, incorporating random misperception (rm) into IDM to model traffic dynamics at intersections with potential perception errors. li_unraveling_2024presented the Dynamic Cooperative IDM (DC-IDM), addressing information errors and their impact on CAVs. liu_connected_2022 proposed a microscopic model for mixed traffic flow unsignalized intersections for autonomous driving, which incorporates the Ornstein–Uhlenbeck process into IDM to model the random perception errors of vehicles at intersections. luo_modeling_2024 proposed the VC-IDM, integrating a generalized force model to analyze platoon behavior and defense mechanisms under cyberattack scenarios.
@@
Summary: The IDM and its extensions owe much of their success to the foundational "intelligent braking strategy," which ensures collision-free behavior when physically possible treiberTrafficFlowDynamics2013. This feature not only enhances IDM’s safety and robustness but also serves as the unifying principle across all deterministic and stochastic variants, making IDM a reliable and flexible framework for both research and practical applications. In deterministic extensions, models like IDM-ACC and HDM refine IDM’s ability to reproduce empirical traffic dynamics, improve stability, and adapt to automated driving scenarios. Meanwhile, stochastic extensions, such as 2D-IDM and SSAM, introduce variability to better capture real-world phenomena like driver heterogeneity, traffic oscillations, and random behavior, thereby enhancing IDM’s realism in diverse traffic conditions.
@@ 3.3.1 Reinforcement Learning-Based Models
Reinforcement learning (RL)-based extensions of IDM aim to improve decision-making and vehicle control by integrating data-driven adaptability with IDM’s physically interpretable structure. For instance, brito_learning_2022 extended IDM using deep reinforcement learning to predict and identify leading vehicles, estimate motion plans, compute control commands (e.g., acceleration), and optimize trajectories in dense traffic, thereby enhancing safety and feasibility. Similarly, caoTrustworthySafetyImprovementAutonomous2022a developed a decision framework for RL, incorporating IDM as a lower performance bound, where autonomous vehicles (AVs) only adopt RL policies if they outperform IDM.
Additionally, baiLongitudinalControlAutomatedVehicles2024a combined Twin Delayed Deep Deterministic Policy Gradient (TD3)-based RL with IDM to achieve improved longitudinal control, safety, and efficiency. kerbel_shared_2024 incorporated IDM’s expected acceleration as a state variable in shared learning to enhance fleet-wide fuel economy and adaptability. Furthermore, hartRobustCarfollowingBasedDeep2024 proposed an RL model with reward functions grounded in IDM’s assumptions, using parameters such as the desired time gap, maximum and comfortable accelerations/decelerations, and desired speed to guide learning.
@@ 3.3.2 Machine and Deep Learning-Based Models
The integration of IDM with machine learning (ML) and deep learning (DL) methods has enabled advanced modeling of vehicle behavior by blending data-driven insights with the established physical frameworks of IDM. For example, chen_investigating_2020 developed an LSTD-IDM model that combines long-term and short-term driving characteristics by merging IDM with longitudinal driving data. moPhysicsInformedDeepLearningParadigm2021c proposed a physics-informed deep learning framework (PIDL-CF), blending IDM with deep learning to predict driver acceleration under various conditions.
Hybrid models have also extended IDM with probabilistic and transformer-based approaches. bhattacharyya_hybrid_2022 introduced a hybrid particle-filtering method combined with collaborative IDM extensions. gengPhysicsInformedTransformerModelVehicle2023b presented the Transformer-IDM (PIT-IDM) for enhanced vehicle trajectory prediction by integrating IDM with deep learning techniques. Similarly, kangTrajectoryBasedEmbeddingRandomCoefficients2023b proposed PGM-IDM, combining trajectory-based embeddings with probabilistic graphical models to improve vehicle-following responses.
Further advancements include liuQuantileRegressionPhysicsInformedDeepLearning2023c, which introduced QRPIDL-CF, fusing quantile regression and deep learning with IDM’s physical insights. tang_grounded_2024 proposed a grounded relational reasoning approach that incorporates IDM’s expected distance into the reward function of a Markov decision process, while zhang_bayesian_2024 developed a memory-augmented Bayesian calibration (MA-IDM) to address parameter uncertainties and behavioral differences in IDM, significantly enhancing its predictive accuracy.
@@ 3.4.1 Deterministic Model Comparisons
Many studies have compared their proposed models with IDM, but relatively few have conducted comprehensive and unified evaluations that fairly compare all models, including IDM. One such study is punzoCalibrationCarFollowingDynamicsAutomated2021, where deterministic models, including IDM, Gipps’ model, and the full velocity difference model (FVDM) with constant-time headway (CTH) and SIGMOID spacing policies, were systematically evaluated across multiple datasets and calibration settings. IDM demonstrated strong and consistent calibration performance, achieving low errors in speed and acceleration while performing well in spacing accuracy. Although Gipps’ model slightly outperformed IDM in reproducing spacing for both human-driven and autonomous vehicle datasets, IDM maintained a more balanced performance across all three measures: spacing, speed, and acceleration. The robustness of IDM across diverse datasets and calibration settings, especially when using Pareto-efficient objective functions such as NRMSE(s,v,a), underscores its versatility as a reliable model for CF dynamics.
@@ 4.1 Applications in Traffic Stability Analysis
A significant focus has been on understanding traffic stability through IDM. kestingHowReactionTimeUpdate2008 initiated this exploration by analyzing the influence of reaction time and speed adaptation on traffic stability, using IDM to reveal various instability mechanisms. Building on this, jinEquivalenceContinuumCarFollowingModels2016a demonstrated the equivalence between continuum and CF models, illustrating that higher-order IDM can be transformed into continuous models with consistent stability conditions. monteil_linear_2014 advanced the discussion by examining string stability in mixed traffic flows and the impact of multiple anticipations in cooperative CF models. montaninoHomogeneousHeterogeneousTrafficFlows2021b further delved into string stability in mixed and heterogeneous traffic flows, incorporating uncertainties in model parameters. They later proposed a unified modeling framework (montaninoStringStabilityMixedHeterogeneous2021b) to assess stability, considering driver and vehicle heterogeneity and using IDM to verify this framework. bouadiStochasticFactorsStringStability2022b extended these concepts into the realm of stochastic models, investigating string stability in stochastic CF models using IDM with stochastic variants and the generalized Lyapunov equation to establish stability conditions.
@@ 4.2 Applications in Calibration Framework
For the development of calibration frameworks, building on this focus on error, jin_reducing_2014 explored error accumulation in CF models calibrated with vehicle trajectory data, deriving a dynamic error model based on acceleration. longBiScaleCarFollowingModelCalibration2024a introduced a bi-scale CF model calibration using corridor-level trajectories, highlighting macro measurement error accumulation. To address the validation aspect, treiberValidationTrafficFlowModels2012 developed a quantitative method for calibrating and validating traffic flow models based on the spatiotemporal evolution of congested traffic patterns. In terms of parameter importance and simplification, punzo_we_2015 conducted sensitivity analysis on driving behavior models, finding leader trajectory more important than parameters, and proposed a simplified model with faster convergence. For setting robust calibration objectives, punzoCalibrationCarFollowingDynamicsAutomated2021 proposed robust guidelines for selecting MoP and GoF for calibrating CF dynamics. They used Pareto efficiency and indifference curves for function comparison. Focusing on data integrity, sharmaMoreAlwaysBetterImpact2019b investigated the impact of vehicular trajectory completeness on the calibration and validation of CF models. Finally, addressing behavioral pattern recognition, zhengMultiObjectiveCalibrationFrameworkCapturing2023b proposed a multi-objective calibration framework to capture behavioral patterns of autonomously-driven vehicles, validating with empirical observations of ACC vehicle dynamics. zhouCalibrationStochasticCarFollowing2025a established the calibration framework of stochastic CF models and revealed the essential differences between deterministic and stochastic CF model calibration using various stochastic IDM variants.
@@
As for the development of the optimization algorithms for calibration, rahman_improving_2015 proposed a stochastic calibration method using Bayesian estimation and MCMC simulation. liGlobalOptimizationAlgorithmTrajectory2016b developed a global optimization algorithm combining direct and gradient searches for CF models. zhongCrossEntropyMethodProbabilisticSensitivity2016b introduced a calibration method using the Cross-Entropy Method and sensitivity analysis under noisy data. fardCopulaBasedEstimationDistributionAlgorithm2019b suggested a copula-based algorithm to explore complex parameter interactions in traffic models. wangIntegratedSelfConsistentMacroMicroTraffic2024a created an integrated macro-micro calibration framework using enhanced optimization and deep learning.
@@
In addition to the above theme, calibration is crucial for model evaluation. zhuModelingCarFollowingBehaviorUrban2018b investigated the applicability of CF models for Chinese drivers, highlighting IDM’s effectiveness and portability on Shanghai highways. campi_roundabouts_2024 assessed various models in roundabout scenarios, with IDM excelling in capturing CAV behavior. zhouExperimentalFeaturesEmissionsFuel2023b used four classical emission and fuel consumption models and found IDM effective at the vehicle pair level for energy consumption and emissions predictions. However, at the platoon level, predicted results were almost constant, contrasting with experimental observations.
@@
Based on calibrated IDM, different behavior patterns and vehicle types can be extracted. ossenHeterogeneityCarFollowingBehaviorTheory2011b used data collected by helicopter to calibrate the IDM to identify differences in following behaviors between different driver types, showing significant heterogeneity. makridis_response_2020 estimated the response time of vehicles equipped with ACC by calibrating IDM parameters and found that it was comparable to that of human drivers. hu_autonomous_2023 calibrated IDM parameters to distinguish the empirical following behaviors of AVs and HVs in Waymo data. li_automated_2023 proposed an AV identification algorithm based on IDM calibration results, that is, using error results to distinguish AVs from HVs.
@@
Summary: IDM has played a pivotal role in the development of calibration frameworks, supporting advances in model performance metrics, optimization algorithms, and behavioral feature extraction. From error modeling to cross-scenario validation, IDM has been widely applied due to its analytical tractability and interpretability. However, most calibration efforts still focus on deterministic model variants, while the calibration of stochastic IDM-family models remains notably underdeveloped due to challenges in parameter identifiability, noise disentanglement, and the lack of robust likelihood functions (punzoCalibrationCarFollowingDynamicsAutomated2021). This limits their empirical applicability despite their theoretical promise. Moving forward, calibration should be re-conceptualized not merely as parameter fitting but as a behavioral inference task, requiring methods that can accommodate uncertainty, heterogeneity, and model adaptivity across varying traffic regimes.